# Unveiling AI’s Blind Spots: An Oracle for In-Domain, Out-of-Domain, and Adversarial Errors

### v1

@@ 5 Conclusion
S1 Detailed Performance of Mentors across various error sources
@@
As mentioned in Sec. 4.1, the detailed results of mentors across various error sources for the CIFAR-10, CIFAR-100, ImageNet-1K datasets are shown in Fig. S1, Fig. S2 and Fig. S3 respectively.
S2 Detailed performance of mentors across mentee architectures
@@
In Fig. 5, we show the generalization performance of mentors averaged over all three error types of mentees with various architectures. Here, we expand the results in the form of tables listing out all the individual accuracy for all the error sources on CIFAR-10, CIFAR-100 and ImageNet-1K datasets in Tab. S1, Tab. S2, and Tab. S3 respectively.
@@
S3 Detailed results of the ablation study on the loss components in SuperMentor
Extending the results shown in Tab. 2, we now include their standard deviations after 3 runs as presented in Tab. S4.
| OOD | AA | |||||||||||
| ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |||

### v2

@@ Appendix I SuperMentor performance on additional OOD domains on ImageNet dataset
Appendix K Detailed performance of mentors across various error sources
@@
As mentioned in the caption of Fig. 3, the detailed results of mentors across various error sources for the CIFAR-10, CIFAR-100, ImageNet-1K datasets are shown in Appendix, Fig. S3, Fig. S4 and Fig. S5 respectively.
Appendix L Detailed performance of mentors across mentee architectures
@@
In Fig. 5, we show the generalization performance of mentors averaged over all three error types of mentees with various architectures. Here, we expand the results in the form of tables listing out all the individual accuracy for all the error sources on CIFAR-10, CIFAR-100 and ImageNet-1K datasets in Appendix, Tab. S6, Tab. S7, and Tab. S8 respectively.